OpenStudy (anonymous):

Can someone check my answer? I just want to make sure I have this right (: Part 1: Use the quadratic formula to solve x2 + 8x = –2. (3 points) Part 2: Using complete sentences, explain the process you used. (2 points) Part 3: Why is the quadratic formula the best method to use? (2 points)

OpenStudy (anonymous):

Here is my final answer: -4 +/- sqrt 14 over 2

OpenStudy (anonymous):

x = -4 +/- sqrt 14 over 2 (forgot the "x =")

OpenStudy (anonymous):

Sweet! Thanks for checking :D

OpenStudy (anonymous):

Can you help me out with part 3? I want to say, "because it is almost always reliable" but I don't think that'll work lol

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

There's a slight mistake here..hold on.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[x ^{2}+8x + 2 =0\] \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] \[\frac{ -8 \pm \sqrt{8^{2}-4(1)(2)} }{ 2 }\] \[\frac{ -8 \pm \sqrt{64 - 8} }{ 2 }\] \[\frac{ -8 \pm \sqrt{56} }{ 2 }\] \[\frac{ -8 \pm 2\sqrt{14} }{ 2 }\] \[\frac{ 2(-4 \pm \sqrt{14}) }{ 2 }\] \[-4 \pm \sqrt{14}\]

OpenStudy (anonymous):

I see what I did wrong. I divided the bottom 2 when I should have just left it alone

OpenStudy (anonymous):

Part 3: The answer here is preferential, but I feel that: The quadratic formula can be used to find the roots for any quadratic equation, and as such, is more reliable and faster than either completing the square or trying to factor our the solutions.

OpenStudy (anonymous):

Okay, thanks! You were extremely helpful :D

OpenStudy (anonymous):

It's fine.